A Carnot cycle has a thermal efficiency of 32%. If the work developed is 1000 watt-hour, determine the heat supplied.


1
Expert's answer
2021-11-22T17:51:25-0500

Given quantities:

"\eta = 0.32 \space \space \space A_1=1000 \space \space \space A_2-?"

"\eta = \large\frac{A_1-A_2}{A_1} \to" "A_2 = A_1(1-\eta)=1000(1-0.32)=680"

Answer: 680 watt-hour